OpenStudy (anonymous):

A bike shop sells rolling strollers for runners with babies, with a total of 28 in the store. Some of the strollers have 3 wheels and some have 4. If there are 100 wheels, how many 3-wheel strollers are there? A. 25 B. 12 C. 3 D. 16 Please help

OpenStudy (anonymous):

3X+4Y=100 X+Y=2

OpenStudy (anonymous):

X+Y=28

OpenStudy (math&ing001):

Let x be the number of 3-wheel strollers and y the number of 4-wheel strollers 28=x+y 3x+4y=100 Solve the system for x and y.

OpenStudy (anonymous):

that would make it 25?

OpenStudy (anonymous):

no that would actually make it D 16 right?

OpenStudy (anonymous):

sorry 16 yes

OpenStudy (anonymous):

awesome thanks @harpreetsk and thank you too @math&ing001

OpenStudy (math&ing001):

Actually 16 is the number of 4-wheell strollers so 12 the answer :)

OpenStudy (anonymous):

Ohhhh yeah that makes more sense because 16 is a multiple of 4 and 12 is a multiple of 3

OpenStudy (anonymous):

thanks @math&ing001

OpenStudy (math&ing001):

Welcome